OpenStudy (anonymous):

without derivative find the maximum and minimum values of 3sinx+4cosx

OpenStudy (anonymous):

This key to doing this question is to know that the sum of a sin function and a cos function is still a sin (or cos) function. Have you come across the equation \[a \sin x + b\cos x = A\sin(x+\theta) \]where \(A=\sqrt{a^2+b^2}\) and \(\tan \theta = \frac{a}{b}\)?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

I expect it's what you have to use. Do you understand what the formula above is saying?